# Beyond RMSE and MAE: Introducing EAUC to unmask hidden bias and unfairness in dyadic regression models

### v1

@@ 5.2 Models
This work employs a diverse array of state-of-the-art dyadic regression models to assess their underlying eccentricity biases and the behaviour of EAUC in measuring them. The models used in this study are Bayesian SVD++ rendle2019difficulty (matrix decomposition), Glocal-K han2021glocal (auto-encoders and convolution kernels), GC-MC berg2017graph (graph auto-encoders), and MF rendle2020neural (simple matrix factorization). We also use two naive predictors as baselines: a random uniform predictor (Random) and a predictor which always outputs the $DMV_{ui}$ of the sample (Dyad Average). This model selection covers different algorithmic approaches used to obtain baseline and state-of-the-art performances in dyadic regression, allowing for a comprehensive evaluation of the prevalence of eccentricity bias, the limitations of global error metrics like RMSE and MAE, and the ability of EAUC to quantify its presence or absence.
5.3 Experimental Implementation Details
This section covers implementation details of models, datasets, and the custom evaluation framework:
  • 1.
@@ 5.3 Experimental Implementation Details
    For each model and dataset from Sections 5.2 and 5.1, we acquire results using the hyperparameters for that dataset described in the model’s original publication, to respect result reproducibility. If such configurations cannot be found in the article, we fall back on those defined in the official code provided by the authors. The full relation of used hyperparameters for each model and dataset can be found in A.
  • 2.
@@
    Model training and evaluation is done on a dedicated machine equipped with 16GB RAM, an NVIDIA GeForce RTX 2060 Super GPU, and an Intel Core i7-10700K CPU @ 3.80GHz.
  • 3.
@@
6 Results and Discussion
The results of our experiments are presented in Table 3, including the performance metrics for each model and dataset. Each entry represents the RMSE, MAE and EAUC for the respective dataset and model. We report results for the state-of-the-art models and the two naive Random and Dyad Average baselines.
Table 3: Global error (RMSE and MAE) and eccentricity bias (EAUC) of the studied models and datasets. The “fairness difficulty” ($D_{KS_{\mathcal{D}}}$) of each dataset is also provided. In each dataset, the first group of results are naive baselines, and the second group are state-of-the-art models. Results represent performance over five runs. For models other than Matrix Factorization (MF), results include only the datasets present in the respective model’s publications. For each dataset and metric, bold and underlined values represent the best and second best-performing model, respectively.
@@ 6 Results and Discussion
The results reveal that, when comparing performances between different datasets, a lower RMSE does not correlate with a lower bias (and therefore a lower EAUC). For instance, models trained on Movielens 1M showcase lower RMSE than those trained on Movielens 100K, yet exhibit equally high EAUC values and therefore similarly severe eccentricity bias. Because the optimization criterion is similar for all the models evaluated (minimizing mean errors), if we consider a single dataset we observe models with a slightly better RMSE exhibit a still high but slightly lower EAUC; in Section 6.1, we demonstrate that this slight difference between high EAUC values is not sufficient to consider the lower better model as sufficiently unbiased. Table 3 also contains the $D_{KS_{\mathcal{D}}}$ for each dataset: it can be observed that datasets with a higher $D_{KS_{\mathcal{D}}}$ have higher EAUC for the trained models, indicating that less uniform local distributions of observed values of users and items lead to learnt models with more severe eccentricity biases.
@@
When eccentric cases hold significant importance, such as in individualized drug dosage prediction, avoiding large errors in eccentric cases can be more critical for system usability and fairness than seeking near-zero errors for common, predictable scenarios. Evaluating such tasks should not be limited to global error metrics like RMSE or MAE, but also measuring entity-wise biases, which we achieve with the proposed EAUC, to promote model fairness.
@@
Figure 4: Relationship between the eccentricity of observed values and the prediction error (lower is better) for each model and dataset. Mean and standard deviation of five runs are reported. For models other than MF, results are reported for datasets used in each model’s original publication.
6.1 Correlating lower EAUC and lighter eccentricity bias
@@
As discussed in Section 4.2, an unbiased model should prioritize sufficiently good predictions no matter the underlying eccentricity of the sample, rather than excelling only at predicting the $DMV_{ui}$ of the $(u,i)$ dyad -by definition, the most common test case- but making worse-than-random predictions for deviated samples with equal importance in real-world tasks. To illustrate this, and to demonstrate the ability of EAUC to discern the degree of eccentricity bias a model suffers from (a lower EAUC directly correlating with a less biased model, and vice-versa), we experimented with simple linear corrections as naive prediction de-biasing techniques.
@@
For all datasets, the Train set is separated into the final Train set, used to train the dyadic regression model, and a small “Correction” set, used to compute the necessary de-biasing corrections based on the predictive behaviour observed on the trained model. In the Correction set, to uniformize the observed values distributions of users and items, we segregated the samples $(u,i,r_{ui})$ in 10x10 equally spaced bins depending on their user’s and item’s average observed values $\bar{r}_{u}$ and $\bar{r}_{i}$, and applied ML-RUS undersampling charte2015addressing to each bin to obtain the same number of samples per unique $r_{ui}$ value and bin: this avoids the learnt correction from succumbing from the same inductive bias as the original model. The correction function is then obtained by fitting on the undersampled Correction set a linear regression that takes as inputs the -possibly biased- value $\hat{r}_{ui}$ predicted by MF and the user’s and item’s average observed values $\bar{r}_{u}$ and $\bar{r}_{i}$, and must output $r_{ui}$, the real -unbiased- observed value of the sample, such that $\hat{r}_{ui_{\text{unbiased}}}=f_{corr}(\hat{r}_{ui_{\text{biased}}},\bar{r}_{u},\bar{r}_{i})$. In essence, this correction tries to learn the degree of eccentricity bias and towards which average observed value the predictions are skewed depending on the past observed values of the user and item, to appropriately de-bias Test set predictions by approximating them to their likely real value.
@@
Table 4 and Figure 5 show the effects of the corrections in the eccentricity bias and the different metrics when applied to an MF model and evaluated in each dataset’s Test set. We use two different approaches to rescale the corrected predictions back to the $[min(r_{ui}),max(r_{ui})]$ range: the Clipping approach clips the estimations to said range, while the Sigmoid approach applies a sigmoidal function and then rescales the output to the range, to avoid a collapse of corrected predictions toward the extreme values. Although they should not be interpreted as a foolproof de-biasing prediction, their effects on the prediction behaviour of MF are evident. While the corrections perform worse than the trained models for near-zero eccentricities, they maintain a much more restrained error as eccentricity increases. Importantly, this de-biasing correction exhibits a correspondence with a lower EAUC across all datasets, indicating that a conscious mitigation of the bias does signify a reduction in EAUC and vice-versa.
Table 4: Global error (RMSE and MAE) and eccentricity bias (EAUC) of an MF model and the two alternative de-biasing techniques through linear corrections (estimation of the real value from the biased prediction) with Clipping or Sigmoid rescaling. For each dataset and metric, bold and underlined values represent the best and second best-performing model, respectively.
@@
7 Conclusions and Future Directions
This study provides a comprehensive examination of the impact of non-uniform local user and item observed value distributions on regression models over dyadic data. We uncover a consistent bias across state-of-the-art models and datasets, where predictions tend to severely gravitate towards user and item average observed values, leading to worsened errors in more eccentric samples, even surpassing the errors of a uniform random prediction. While RMSE and MAE, used ubiquitously as the sole evaluation metrics in the task, offer global insights into the prediction error of models, they fall short in capturing biases such as the aforementioned, posing risks of unfairness and potentially hazardous implications depending on the system’s final application.
@@ 7 Conclusions and Future Directions
8 Acknowledgements
This work has been supported by the National Plan for Scientific and Technical Research and Innovation of the Spanish Government (Grant PID2019-109238GB-C22), by the Spanish Ministry of Science and Innovation (Grant FPU21/05783), and by the Xunta de Galicia (Grant ED431C 2022/44) with the European Union ERDF funds. CITIC, as Research Center accredited by Galician University System, is funded by “Consellería de Cultura, Educación e Universidade from Xunta de Galicia”, supported in an 80% through ERDF Operational Programme Galicia 2014-2020, and the remaining 20% by “Secretaría Xeral de Universidades” (Grant ED431G 2019/01).
@@ Appendix A Model Hyperparameters
Table 5 below summarizes the hyperparameter configurations employed for various models across different datasets in the experimental setup. These configurations were either adopted from the original publications or extracted from the official repositories and source codes.
Table 5: Hyperparameter configurations for the models and datasets considered in this work.

### v2

@@ V-B Models
This work employs a diverse array of state-of-the-art dyadic regression models to assess their underlying eccentricity biases and the behaviour of EAUC in measuring them. The models used in this study are Bayesian SVD++ [24] (matrix decomposition), Glocal-K [25] (auto-encoders and convolution kernels), GC-MC [26] (graph auto-encoders), and MF [17] (simple matrix factorization). We also use two naive predictors as baselines: a random uniform predictor (Random) and a predictor which always outputs the $DMV_{ui}$ of the sample (Dyad Average). This model selection covers different algorithmic approaches used to obtain baseline and state-of-the-art performances in dyadic regression, allowing for a comprehensive evaluation of the prevalence of eccentricity bias, the limitations of global error metrics like RMSE and MAE, and the ability of EAUC to quantify its presence or absence.
V-C Experimental Implementation Details
@@
For each model and dataset, we acquire results using the hyperparameters for that dataset described in the model’s original publication, to respect result reproducibility. If such configurations cannot be found in the article, we fall back on those defined in the official code provided by the authors. The full relation of used hyperparameters for each model and dataset can be found in Appendix A.
@@
Model training and evaluation is done on a dedicated machine with 16GB RAM, an NVIDIA GeForce RTX 2060 Super GPU, and an Intel Core i7-10700K CPU @ 3.80GHz.
@@
VI Results and Discussion
The results Table III includes the performance metrics (RMSE, MAE, EAUC) obtained for each model and dataset. We report results for the state-of-the-art models and the two naive Random and Dyad Average baselines.
TABLE III: Global error RMSE and MAE and eccentricity bias EAUC per model and dataset, and “fairness difficulty” $D_{KS_{\mathcal{D}}}$ per dataset.
@@ VI Results and Discussion
  • In each dataset, the first group of results are naive baselines, and the second group are state-of-the-art models. Results represent performance over five runs. Other than Matrix Factorization (MF), model results include only the datasets present in their respective publications. For each dataset and metric, bold and underlined values represent the best and second best-performing model, respectively.
The results reveal that, when comparing performances between different datasets, a lower RMSE does not correlate with a lower bias (and therefore a lower EAUC). For instance, models trained on Movielens 1M showcase lower RMSE than those trained on Movielens 100K, yet exhibit equally high EAUC values and therefore similarly severe eccentricity bias. Because the optimization criterion is similar for all the models evaluated (minimizing mean errors), if we consider a single dataset we observe models with a slightly better RMSE exhibit a still high but slightly lower EAUC; in Section VI-A, we demonstrate that this slight difference between high EAUC values is not sufficient to consider the lower better model as sufficiently unbiased. Table III also contains the $D_{KS_{\mathcal{D}}}$ for each dataset: it can be observed that datasets with a higher $D_{KS_{\mathcal{D}}}$ have higher EAUC for the trained models, indicating that less uniform local distributions of observed values of users and items lead to learnt models with more severe eccentricity biases.
@@
When eccentric cases hold significant importance, such as in individualized drug dosage prediction, avoiding large errors in eccentric cases is more critical for system usability and fairness than seeking near-zero errors in common, predictable scenarios. Evaluating such tasks should not be limited to global error metrics like RMSE or MAE, but also measuring entity-wise biases, which we achieve with the proposed EAUC, to promote model fairness.
@@
Figure 4: Relationship between eccentricity of observed values and prediction error (lower is better) for each model and dataset (mean and standard deviation of five runs). For models other than MF, results are reported for datasets used in each model’s original publication.
VI-A Correlating lower EAUC and lighter eccentricity bias
@@
As previously discussed, an unbiased model should prioritize sufficiently good predictions no matter the underlying eccentricity of the sample, rather than excelling only at predicting the $DMV_{ui}$ of the $(u,i)$ dyad -by definition, the most common test case- but making worse-than-random predictions for deviated examples with equal importance in real-world tasks. To demonstrate the ability of EAUC to measure the severity of a model’s eccentricity bias (a lower EAUC directly correlating with a less biased model, and vice-versa), we experimented with simple linear corrections as naive prediction de-biasing techniques.
@@
For all datasets, the Train set is separated into the final Train set, used to train the dyadic regression model, and a small “Correction” set, used to compute the necessary de-biasing corrections based on the predictive behaviour observed on the trained model. In the Correction set, to uniformize the observed values distributions of users and items, we segregated the samples $(u,i,r_{ui})$ in 10x10 equally spaced bins depending on their user’s and item’s average observed values $\bar{r}_{u}$ and $\bar{r}_{i}$, and applied ML-RUS undersampling [27] to each bin to obtain the same number of samples per unique $r_{ui}$ value and bin: this avoids the learnt correction from succumbing from the same inductive bias as the original model. The correction function is then obtained by fitting on the undersampled Correction set a linear regression that takes as inputs the -possibly biased- value $\hat{r}_{ui}$ predicted by MF and the user’s and item’s average observed values $\bar{r}_{u}$ and $\bar{r}_{i}$, and must output $r_{ui}$, the real -unbiased- observed value of the sample, such that $\hat{r}_{ui_{\text{unbiased}}}=f_{corr}(\hat{r}_{ui_{\text{biased}}},\bar{r}_{u},\bar{r}_{i})$. In essence, this correction tries to learn the degree of eccentricity bias and towards which average observed value the predictions are skewed depending on the past observed values of the user and item, to appropriately de-bias Test set predictions by approximating them to their likely real value.
@@
Table IV and Figure 5 show the effects of the corrections in the eccentricity bias and the different metrics when applied to an MF model and evaluated in each dataset’s Test set. We use two different approaches to rescale the corrected predictions back to the $[min(r_{ui}),max(r_{ui})]$ range: the Clipping approach clips the estimations to said range, while the Sigmoid approach applies a sigmoidal function and then rescales the output to the range, to avoid a collapse of corrected predictions toward the extreme values. Although they should not be interpreted as a foolproof de-biasing prediction, their effects on the prediction behaviour of MF are evident. While the corrections perform worse than the trained models for near-zero eccentricities, they maintain a much more restrained error as eccentricity increases. Importantly, this de-biasing correction exhibits a correspondence with a lower EAUC across all datasets, indicating that a conscious mitigation of the bias does signify a reduction in EAUC and vice-versa.
TABLE IV: Global error RMSE and MAE and eccentricity bias EAUC of MF and naive de-biasing techniques.
@@
  • The two naive de-biasing techniques employ linear corrections (estimation of the real value from the biased prediction) with Clipping or Sigmoid rescaling. For each dataset and metric, bold and underlined values represent the best and second best-performing model, respectively.
@@
VII Conclusions and Future Directions
This study provides a comprehensive examination of the impact of non-uniform local user and item observed value distributions on regression models over dyadic data. We uncover a consistent bias across state-of-the-art models where predictions tend to severely gravitate towards user and item average observed values, leading to worse-than-random errors in more eccentric samples. While RMSE and MAE, used ubiquitously as the sole evaluation metrics in the task, only measure global prediction errors, they fall short in capturing biases such as the aforementioned, posing risks of unfairness and potentially hazardous implications depending on the system’s final application.
@@ Appendix A Model Hyperparameters
Table V below summarizes the hyperparameter configurations employed for various models across different datasets in the experimental setup. These configurations were either adopted from the original publications or extracted from the official repositories and source codes.
TABLE V: Hyperparameter configurations per model and dataset.